OpenStudy (anonymous):

What is the gravitational force between two identical bodies with a mass of 2.0*104 kg each if they are 3.0 m apart? A. 3.0*10-3 N B. 8.9*10-3 N C. 1.5*10-7 N D. 4.4*10-7 N

OpenStudy (anonymous):

u solved the other one.. u should be able to solve that one too :O

OpenStudy (anonymous):

Well I think its D. But I dont think its correct. @Mashy

OpenStudy (anonymous):

why do you think so? would u mind showing me ur work?

OpenStudy (anonymous):

F = Gm²/d² 3 = Gm²/ 3² 8/G = m² and (G = 6.674281 x 10^(-11)) and then I get lost

OpenStudy (anonymous):

how did u get 8 ? :O

OpenStudy (anonymous):

gravity?

OpenStudy (anonymous):

u had 3/G = m^2/3^2 from there.. how did you get 8/G = m^2 ?

OpenStudy (anonymous):

I have no idea I was working from the book I have no idea how to do it

OpenStudy (anonymous):

u don't know how to solve for m ? from that equation? :-/ that's basic algebra girl! try it!

OpenStudy (anonymous):

I dont understand physics at all and I was sick so I didnt get the lesson taught to me and none of my friends get it either. So if you could just help me out that would be great.

OpenStudy (anonymous):

so u basically have no clue how to even start solving this problem? huh.. ok.. do you know what the universal law of gravitation is? composed by Newton?

OpenStudy (anonymous):

Every point mass in the universe attracts every other point mass with a force that is directly proportional to the product of their masses.

OpenStudy (anonymous):

and? there is something else

OpenStudy (anonymous):

Thats all it says about it in my book...

OpenStudy (anonymous):

are u sure? it says nothing about the distance? at all?

OpenStudy (anonymous):

Nothing at all

OpenStudy (anonymous):

dang.. maybe they have mentioned it in another part or something .. anyhoo.. its inversely proportional to distance square.. that makes sense right? as the objects go faaaar away, the force between them goes down too.. !

OpenStudy (anonymous):

I keep getting 6*10^4

OpenStudy (anonymous):

no.. that's way too high.. and first of all.. i read the question wrong to being with.. gravitational force is asked :D so u have your m1 and m2 which are equal.. you have your d .. and you have your G F = Gm1m2/d^2.. just plug in and see what you end up with?

OpenStudy (anonymous):

G is 2 and M is 3?

OpenStudy (anonymous):

no.. d = distance.. G = universal gravitational constant.. i think you need to spend sometime with your textbook.. don't just try to get an answer for a problem!!..

OpenStudy (anonymous):

Well then how does the 2 and 3 get plugged in if M1 and M2 are equal

OpenStudy (anonymous):

m1 and m2 are the masses.. and read the question.. what are the masses ?

OpenStudy (anonymous):

So d is 3....

OpenStudy (anonymous):

So F = G*2*2/3^2

OpenStudy (anonymous):

so the answer wouldve been A